OpenStudy (anonymous):

What values for theta(0_

OpenStudy (anonymous):

3sin theta = sin theta - 1 or 2sin theta = -1 or sin theta = -1/2 or sin theta = sin -pi/6 because sin -pi/6 is equal to -1/2 or theta = -pi/6